Question:
suppose that the probability that a student will pass a test is four-fifths, and the the students taking the test are all interdependent of one another. what is the probability that none of the four students will pass the test?
Answer: .00160

Answer by dfrazzetto(283)


if p (passing) = 4/5 for each student then p (not passing) = 1-4/5 = 1/5
Since all events are independent, then:
p (4 students not passing) = (1/5)^4 = (1/5)(1/5)(1/5)(1/5) = 0.0016